# Application of Laplace filters to the analysis of lattice time correlators

### v1

@@ Acknowledgements.
The authors would like to gratefully thank to our collaborators within the RBC/UKQCD collaboration, in particular Felix Erben and Andreas Jüttner, for the use of the ensembles and the data presented in this study. AP received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreements No 757646 & 813942. AP is additionally funded in part by UK STFC grant ST/X000494/1. AP was additionally supported by a CERN Scientific Associate position, as well as by a Long Term Invitational Fellowship from the Japan Society for the Promotion of Science. This work used the DiRAC Extreme Scaling service (Tursa / Tesseract) at the University of Edinburgh, managed by the EPCC on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Edinburgh was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure.

### v2

@@ VII Discussion and conclusion
We reiterate that Laplace filtering can be applied in conjunction with the various methods discussed in the Introduction should the resulting correlation matrix not yet be sufficiently well conditioned.
@@ Acknowledgements.
The authors would like to gratefully thank to our collaborators within the RBC/UKQCD collaboration, in particular Felix Erben and Andreas Jüttner, for the use of the ensembles and the data presented in this study. AP received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreements No 757646 & 813942. AP is additionally funded in part by UK STFC grant ST/X000494/1. AP was additionally supported by a CERN Scientific Associate position, as well as by a Long Term Invitational Fellowship from the Japan Society for the Promotion of Science. This work used the DiRAC Extreme Scaling service (Tursa / Tesseract) at the University of Edinburgh, managed by the EPCC on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Edinburgh was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. Two-point correlators used in this work are publicly available at [34], three-point functions are still part of ongoing projects and are available upon reasonable request from the authors.